OpenStudy (anonymous):

Factor: (y - 4)^3 - 125

OpenStudy (campbell_st):

ok this is the difference of 2 cubes \[(y - 4)^3 - 5^3\] this may help

OpenStudy (campbell_st):

the general form is \[a^3 - b^3 = (a - b)(a^2 + ab + b^2)\] so use this for your solution tHope this help.

OpenStudy (anonymous):

\[(y-4)^3-125=((y-4)-5)((y-4)^2+5(y-4)+25)\\=(y-9)(y^2-8y+16+5y-20+25)=(y-4)(y^2-3y+21)\]